Give the 9th term in the sequence 2, -6, 18, -54?
Answer by nerdybill(5404)


What you have is a "geometric progression". Where the following progression holds:
.
An = A1(r^(n-1))
where
An = the "n"th term in the progression
A1 = value of first term
.
See the following web site for more detailed information:
http://en.wikipedia.org/wiki/Geometric_progression
.
In your problem:
- you are asked to find the 9th term (An)
- A1 is 2
- r is -3
.
Plug what we know intoL
An = A1(r^(n-1))
An = 2(-3)^(9-1)
An = 2(-3)^8
An = 2(6561)
An = 13122